If a + b + c = 2024 and 1/a + 1/b + 1/c = 1/2024
find 1/a²⁰²⁵ + 1/b²⁰²⁵ +1/c²⁰²⁵

Answer by mccravyedwin(406)

If a + b + c = 2024 and 1/a + 1/b + 1/c = 1/2024
find 1/a²⁰²⁵ + 1/b²⁰²⁵ +1/c²⁰²⁵
-----------------------------------------------------
Just pick one of the variables to be 2024 and the other two to be
any opposite non-zero numbers, say, k and -k

Say a = 2024, b = k, c = -k

Then

2014 + k - k = 2024 and 1/2024 + 1/k + 1/(-k) = 1/2024

Then

1/2024²⁰²⁵ + 1/k²⁰²⁵ - 1/k²⁰²⁵ = 1/2024²⁰²⁵
  
If you want a decimal approximation and you put it in your
calculator like it is you'll get an overflow because no
ordinary device can handle numbers as large as 2024²⁰²⁵
or as small as 1/2024²⁰²⁵

So here's what you do to get it in scientific notation:

Use base 10 logs:

n = 2024^(-2025)

log(n) = log(2025^(-2025)

log(n) = -2025*log(2024)

log(n) = -2025(3.30621058)

log(n) = -6695.076279

n = 10^(-6695.076279)

n = 10^(-6695)*10^(-0.076279)

n = 10^(-6695)*0.8389208728

Divide the 10^(-6695) by 10 and multiply the 0.8389208728 by 10
 
n = 10^(-6696)*8.389208728

approximately 8.3892078728... × 10^-6696 in scientific notation.

Edwin